solve each equation by completing the square.
n^2+12n+37=2



k^2+8k-47=7

Answer by lwsshak3(11628)

solve each equation by completing the square.
n^2+12n+37=2
(n^2+12n+36)-36+37=2
(n+6)^2=1
n+6=±√1=±1
n=-6±1
n=-7
or
n=-5
..
k^2+8k-47=7
(k^2+8k+16)-16-47=7
(k+4)^2=70
k+4=±√70
k=-4±√70
k=-4-√70≈-12.37
or
k=-4+√70≈4.37